OpenStudy (anonymous):

what is the value of the discriminant -6x^2-4x-13=0

OpenStudy (amistre64):

\[b^2-4(a)(c)\]

OpenStudy (amistre64):

methinks its gone be -

OpenStudy (anonymous):

\[Its -296\]

OpenStudy (jit4won):

D= (-4)^2 -4 (-6)(-13) = 16-312=-296

OpenStudy (amistre64):

i got -292 in me head....

OpenStudy (anonymous):

-296

OpenStudy (amistre64):

i have to interpolate for my inherent stupidity.... so after adjusting, I get -297 lol

OpenStudy (anonymous):

\[I \ think\ it's\ -296\]

OpenStudy (anonymous):

i also got -296. i just like to come here and double check

OpenStudy (anonymous):

\[is \ that\ right\]

OpenStudy (anonymous):

\[become\ a\ fan\ plz\]